OpenStudy (anonymous):

solve the equation -2x=3y=2

OpenStudy (anonymous):

sorry -2x+3y=2

OpenStudy (anonymous):

lol, that makes more sense ;D one sec

OpenStudy (anonymous):

to find "X" intercepts, set Y=0 and solve for "X" -2x+3(0)=2 -2x=2 x=-1 to find "Y" intercepts, set X=0 and solve for "Y" -2(0)+3y=2 3y=2 y=2/3